# Understanding Dataset Distillation via Spectral Filtering

### v1

@@ Abstract
Dataset distillation (DD) has emerged as a promising approach to compress datasets and speed up model training. However, the underlying connections among various DD methods remain largely unexplored. In this paper, we introduce UniDD, a spectral filtering framework that unifies diverse DD objectives. UniDD interprets each DD objective as a specific filter function that affects the eigenvalues of the feature-feature correlation (FFC) matrix and modulates the frequency components of the feature-label correlation (FLC) matrix. In this way, UniDD reveals that the essence of DD fundamentally lies in matching frequency-specific features. Moreover, according to the filter behaviors, we classify existing methods into low-frequency matching and high-frequency matching, encoding global texture and local details, respectively. However, existing methods rely on fixed filter functions throughout distillation, which cannot capture the low- and high-frequency information simultaneously. To address this limitation, we further propose Curriculum Frequency Matching (CFM), which gradually adjusts the filter parameter to cover both low- and high-frequency information of the FFC and FLC matrices. Extensive experiments on small-scale datasets, such as CIFAR-10/100, and large-scale datasets, including ImageNet-1K, demonstrate the superior performance of CFM over existing baselines and validate the practicality of UniDD.
@@ 1 Introduction
The exponential growth of data presents significant challenges to the efficiency and scalability of training deep neural networks [ImageNet, CLIP, ViT, SAM]. Dataset distillation (DD) has emerged as a solution to these challenges, aiming to condense large-scale real datasets into compact, synthetic ones without compromising model performance [DD, DD_survey1, DD_survey2, DD_survey3, DD_survey4]. This approach has shown promise across a range of domains, including image [DC, DM, MTT], time series [TS1, TS2], and graph [GCOND, SGDD, GDEM].
@@
Existing DD methods vary in their optimization objectives, which can be grouped into four main categories. Statistical matching [DM, IDM, CAFE] aligns key statistics, such as mean and variance, between real and synthetic datasets. Gradient matching [DC, IDC, DSA] minimizes the direction of model gradients across real and synthetic data during training. Trajectory matching [MTT, TESLA, FTD, DATM] enforces the synthetic data to emulate the update trajectories of real model parameters. Kernel-based approaches [KIP, FRePo, RFAD] employ closed-form representations to bypass the inner-optimization and improve distillation efficiency. Despite differences in their objectives, all methods strive to minimize the discrepancy between real and synthetic datasets from certain perspectives. This raises some key questions: Are these DD methods related? If so, is there a unified framework that can encompass and explain the objectives of various DD methods? These questions are crucial as a unified framework can deepen our understanding of DD, uncover the essence of existing methods, and offer potential insights for advanced distillation technologies.
@@
As the first contribution of our work, we theoretically analyze some representative DD methods, including DM [DM], DC [DC], MTT [MTT], and FrePo [FRePo], and summarize them into a spectral filtering framework, termed UniDD, uncovering their commonalities and differences. Specifically, these DD methods are all proven to match the feature-feature correlation (FFC) and feature-label correlation (FLC) matrices between the real and synthetic datasets. Their difference lies in the filter function applied to the FFC matrix, which changes the eigenvalues of the FFC matrix to target certain frequency components of the FLC matrix.
Based on the differences in their filtering behaviors, we classify existing DD methods into two categories: low-frequency matching (LFM) and high-frequency matching (HFM). Our experimental investigations, detailed in Section 2.3, demonstrate that LFM-based methods, e.g., DM and DC, capture the global textures, while HFM-based methods, e.g., MTT and FrePo, focus on more fine-grained details. Notably, HFM usually outperforms LFM, suggesting that high-frequency information potentially plays a more significant role in achieving superior distillation performance.
@@
The connections between DD objectives and filter functions further enable us to identify the weaknesses of existing methods. Traditional DD methods typically select the matching objectives based on intuition, while UniDD shifts this paradigm to filter design, making the new DD methods more interpretable and reliable. Based on the guidance of UniDD, we find that existing DD methods only use fixed filter functions during distillation, and can only cover a few frequency components, limiting their adaptability. Therefore, we propose Curriculum Frequency Matching (CFM), which gradually adjusts the filter parameter to increase the ratio of low-frequency information in the high-pass filter, thereby covering a wider range of frequency components. To verify its effectiveness, we conduct extensive experiments on various datasets, including CIFAR-10/100, Tiny-ImageNet, and ImageNet-1k. The results demonstrate that CFM consistently outperforms baselines by substantial margins and yields better cross-architecture generalization ability. The contributions of this paper are summarized below:
@@
  • We introduce UniDD, a spectral filtering framework that interprets each DD objective as a specific filter function applied to FFC and FLC matrices, thus unifying diverse DD methods as a frequency-matching problem.
@@
  • We classify existing methods into low-frequency and high-frequency matching based on their filter functions, highlighting their respective roles in encoding global textures and local details.
@@
Scope of UniDD.
The proposed framework mainly focuses on unifying the objectives of different DD methods, while other research lines, e.g., data parameterization and model augmentation, are beyond the scope of this paper, and we introduce them in the related work section.
2 A Unified Framework of Dataset Distillation
@@
Notations. Let $\mathcal{T}=(H,Y)$ denote a real dataset, where $H$ represents the original data with $|H|=n$ samples, and $Y\in\mathbb{R}^{n\times c}$ is the one-hot label matrix for $c$ classes. The goal of DD is to learn a synthetic network $\mathcal{S}=(H_{s},Y_{s})$, where $|H_{s}|=m\ll n$ and $Y_{s}\in\mathbb{R}^{m\times c}$, such that models trained on $\mathcal{T}$ and $\mathcal{S}$ have comparable performance. In addition, there is a pre-trained distillation network $\phi(\cdot)$, such as ConvNet [DC] or ResNet-18 [ResNet] for the image datasets. We use $X=\phi(H)\in\mathbb{R}^{n\times d}$ and $X_{s}=\phi(H_{s})\in\mathbb{R}^{m\times d}$ to indicate the data representations learned by the distillation network on real and synthetic datasets, where $d$ is the dimension.
Furthermore, we introduce two basic assumptions that have been widely adopted in previous methods [DD_survey1, FRePo] to simplify the analysis of DD:
Assumption 1.
The distillation process assumes a fixed feature extractor, reducing the problem to optimizing a linear classifier $W\in\mathbb{R}^{d\times c}$ with the loss function $\left\|XW-Y\right\|_{F}^{2}$.
Assumption 2.
The training dynamics of the linear classifier are analyzed under full-batch gradient optimizations.
Based on the above preliminaries, we propose UniDD, a spectral filtering framework that unifies a wide range of DD methods, formulated as follows:
where $X^{\top}X,X_{s}^{\top}X_{s}\in\mathbb{R}^{d\times d}$ indicates the FFC matrices of the real and synthetic datasets and $X^{\top}Y,X_{s}^{\top}Y_{s}\in\mathbb{R}^{d\times c}$ denote the FLC matrices. $f(\cdot)$ is the filter function and $g(\cdot)$ is a binary function with $g(X^{\top}Y)=I$ or $X^{\top}Y$.
Spectral Filtering [SFT].
@@
Since $X^{\top}X$ is always positive semi-definite, it can be decomposed into $X^{\top}X=U\Lambda U^{\top}$, where $U$ is the eigenvectors and $\Lambda$ is the diagonal eigenvalue matrix with $\lambda_{i}=\Lambda_{ii}$. Without loss of generality, we assume that the eigenvalues are sorted in a descent order, i.e., $\lambda_{1}\geq\cdots\geq\lambda_{d-1}\geq\lambda_{d}\geq 0$. The filter function $f(\cdot)$ acting on $X^{\top}X$ is equivalent to acting on its eigenvalues:
$f(X^{\top}X)g(X^{\top}Y)=Uf(\Lambda)U^{\top}g(X^{\top}Y),$ (2)
where the FFC and FLC matrices serve as filter and signal, respectively. The matrix $U^{\top}$ is called Fourier transform and is used to convert signals into the spectral domain, and $f(\Lambda)$ is a filter that changes the energy of different frequency components. Depending on the property of $f(\cdot)$, filters can be divided into different categories, such as low-pass and high-pass. For example, a low-pass filter can be defined as $f(\lambda_{i})\geq f(\lambda_{i+1}),\forall i\in[1,d-1]$, as a larger eigenvalue indicates a lower frequency.
We analyze four representative DD methods in the following and explain their relationship with different filter functions. Based on the filter behaviors, they can be divided into LFM-based and HFM-based methods. See Table 1 for a quick overview and Appendix 1 for detailed derivation.
@@
Low-frequency DM [DM] $\left\|X^{\top}Y-X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=1$
DC [DC] $\left\|X^{\top}X-X_{s}^{\top}X_{s}\right\|_{F}^{2}+\left\|X^{\top}Y-X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=\{1,\lambda\}$
High-frequency MTT [MTT] $\left\|(I-\alpha X^{\top}X)^{P}-(I-\alpha X_{s}^{\top}X_{s})^{Q}\right\|_{F}^{2}+$ $\alpha\left\|\sum_{p=0}^{P-1}(I-\alpha X^{\top}X)^{p}X^{\top}Y-\sum_{q=0}^{Q-1}(I-\alpha X_{s}^{\top}X_{s})^{q}X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=(1-\alpha\lambda)^{\{p,q\}}$
FRePo [FRePo] $\left\|(X^{\top}X+\beta I)^{-1}X^{\top}Y-(X_{s}^{\top}X_{s}+\beta I)^{-1}X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=(\lambda+\beta)^{-1}$
@@ 2.1 Low-frequency Matching
Statistical Matching.
Matching important statistics between the real and synthetic datasets is a straightforward way to distill knowledge. DM [DM] proposes to match the average representations of each class, whose objective function can be defined as:
@@
where the corresponding filtering functions are $f(X^{\top}X)=I$ and $g(X^{\top}Y)=X^{\top}Y$, respectively. For clarity, we omit the mean normalization here, but this does not affect the definitions of $f(\cdot)$ and $g(\cdot)$.
However, DM does not perform well on complex distillation backbones, e.g., ResNet-18, as it only matches the statistic of the last layer of backbones. To overcome this limitation, SRe2[SRe2L] proposes to match the mean and variance information in each Batch Normalization (BN) layer. The objective function is defined as:
@@
Gradient Matching.
Another LFM example is gradient matching [DC], which minimizes the differences of model gradients in the real and synthetic datasets. The gradients of the linear classifier are calculated as:
@@ 2.2 High-frequency Matching
Trajectory Matching.
Matching the long-range training trajectories has been proven to be an effective approach for DD. We take MTT [MTT] as an example, which minimizes the differences between network parameters trained on the real and synthetic datasets. The objective function can be formulated as:
@@
where $\alpha$ is the learning rate and $W^{0}$ indicates the initial parameters. After optimizing $K$ iterations, we have:
@@
Subsequent works improve MTT from different perspectives, such as memory overhead [TESLA], trajectory training [FTD], and trajectory selection [DATM]. However, they all follow the same objective function so that they can be naturally included in our unified framework.
KRR.
The early-stage DD algorithms usually update the synthetic data by solving a two-level optimization problem. To reduce computation and memory costs, KRR-based methods are proposed to replace the inner-optimization with a closed-form solution:
@@
By applying a matrix identity transform [welling2013kernel] on the second term, the objective function can be unified into UniDD:
@@
It is worth noting that only the linear kernel case can be unified into our framework, e.g., KIP [KIP] and FrePo [FRePo]. We leave the analysis of the non-linear kernel, e.g., Gaussian and polynomial, as further work.
2.3 Explanation of Different Filters
(a) All-pass, $f(\lambda)=1$
(b) Low-pass, $f(\lambda)=\lambda$
(c) High-pass, $f(\lambda)=(1-0.5\lambda)^{P}$
(d) High-pass, $f(\lambda)=(\lambda+\beta)^{-1}$
Figure 1: Visualization of different filtering functions. A larger eigenvalue indicates a lower frequency. If $f(\lambda)$ has a large value when $\lambda$ is small, then $f(\cdot)$ is a high-pass filter, and vice versa.
(a) Raw image
(b) Low-pass
(c) $\beta=0.1$
(d) $\beta=0.01$
Figure 2: Synthetic images distilled by different filters.
The above analysis emphasizes that the design of filter functions $f(\cdot)$ plays a vital role in DD. Therefore, this section further analyzes the differences between filters, including all-pass, low-pass, and high-pass. See Figure 1 for the visualizations of different filters.
The all-pass filter maps all eigenvalues to 1 and maintains all frequency information of the input signal. Although there is no information loss, the filter will make the synthetic dataset converge to a trivial solution, i.e., $X_{s}^{\top}X_{s}=I$, which cannot encode the knowledge of the real dataset, i.e., $X^{\top}X$.
Low-pass filters amplify the impact of low-frequency information, which represents the global texture of natural images. Similarly, we find that matching the low-frequency information of the FFC matrix enables synthetic images to learn the texture of the real dataset. Figure 2(a) is a raw goldfish image in Tiny-ImageNet. When applying a low-pass filter on the FFC matrix, the synthetic image, shown in Figure 2(b), becomes blurred.
In contrast, high-pass filters emphasize the local fine-grained details of the input data. Figures 2(c) and 2(d) illustrates the synthetic images learned by the high-pass filter $f(\lambda)=(\lambda+\beta)^{-1}$ with different hyper-parameters. The results demonstrate that smaller values of $\beta$ allow the synthesized images to capture more intricate details. Thus, by tuning the value of $\beta$, we can modulate the frequency distribution within the synthetic images.

### v2

@@ Abstract
Dataset distillation (DD) has emerged as a promising approach to compress datasets and speed up model training. However, the underlying connections among various DD methods remain largely unexplored. In this paper, we introduce UniDD, a spectral filtering framework that unifies diverse DD objectives. UniDD interprets each DD objective as a specific filter function applied to the eigenvalues of the feature-feature correlation (FFC) matrix to extract certain frequency information of the feature-label correlation (FLC) matrix. In this way, UniDD reveals that the essence of DD fundamentally lies in matching frequency-specific features. Moreover, we characterize the roles of different filters. For example, low-pass filters, e.g., DM and DC, capture blurred patches, while high-pass filters, e.g., MTT and FrePo, prefer to synthesize fine-grained textures and have better diversity. However, existing methods can only learn the sole frequency information as they rely on fixed filter functions throughout distillation. To address this limitation, we further propose Curriculum Frequency Matching (CFM), which gradually adjusts the filter parameter to cover both low- and high-frequency information of the FFC and FLC matrices. Extensive experiments on small-scale datasets, such as CIFAR-10/100, and large-scale ImageNet-1K, demonstrate the superior performance of CFM over existing baselines and validate the practicality of UniDD.
@@ 1 Introduction
The exponential growth of data presents significant challenges to the efficiency and scalability of training deep neural networks. Dataset distillation (DD) has emerged as a solution to these challenges, aiming to condense large-scale real datasets into compact, synthetic ones without compromising model performance (DD; DD_survey1; DD_survey2; DD_survey3; DD_survey4). This approach has shown promise across a range of domains, including image (DC; DM; MTT), time series (TS1; TS2), and graph (GCOND; SGDD; GDEM).
@@
Existing DD methods vary in their optimization objectives, which can be grouped into four main categories. Statistical matching (DM; IDM) aligns key statistics, such as mean and variance, between real and synthetic datasets. Gradient matching (DC; IDC; DSA) minimizes the direction of model gradients across real and synthetic data during training. Trajectory matching (MTT; DATM) enforces the synthetic data to emulate the update trajectories of real model parameters. Kernel-based approaches (FRePo; RFAD) employ a closed-form solution to bypass the inner optimization and improve distillation efficiency. Despite differences in their objectives, all methods strive to minimize the discrepancy between real and synthetic datasets from certain perspectives. This raises some key questions: Are these DD methods related? If so, is there a unified framework that can encompass and explain the objectives of various DD methods? These questions are crucial as a unified framework can deepen our understanding of DD, uncover the essence of existing methods, and offer potential insights for advanced distillation technologies.
@@
As the first contribution of our work, we theoretically analyze some representative DD methods, including DM (DM), DC (DC), MTT (MTT), and FrePo (FRePo), and summarize them into a spectral filtering framework, termed UniDD, uncovering their commonalities and differences. Specifically, these DD methods are all proven to match the feature-feature correlation (FFC) and feature-label correlation (FLC) matrices between real and synthetic datasets. Their difference lies in the filter function applied to the FFC matrix, which changes the eigenvalues of the FFC matrix to extract certain frequency components of the FLC matrix. Based on their filtering behaviors, we classify existing DD methods into low-frequency matching (LFM) and high-frequency matching (HFM). Our experimental investigations, detailed in Section 4, demonstrate that LFM-based methods, e.g., DM and DC, tend to learn coarse-grained colors and blur the synthetic images. In contrast, HFM-based methods, e.g., MTT and FrePo, focus on fine-grained textures, which have better diversity.
@@
The connections between DD objectives and filter functions further enable us to identify the weaknesses of existing methods. Traditional DD methods typically select the matching objectives based on intuition, while UniDD shifts this paradigm to filter design, making the new DD methods more interpretable and reliable. Based on the guidance of UniDD, we find that existing DD methods only use fixed filter functions during distillation and can only learn a single frequency component, which limits their adaptability. Therefore, we propose Curriculum Frequency Matching (CFM), which gradually adjusts the filter parameter to increase the ratio of high-frequency information in the synthetic data, thereby covering a wider range of frequency components. To verify its effectiveness, we conduct extensive experiments on CIFAR-10/100, Tiny-ImageNet, and ImageNet-1k. The results demonstrate that CFM consistently outperforms baselines by substantial margins and yields better cross-architecture generalization ability. The contributions of this paper are summarized below:
@@
  • We introduce UniDD, a spectral filtering framework that interprets each DD objective as a specific filter function applied to the FFC and FLC matrices, thus unifying diverse DD methods as a frequency-matching problem.
@@
  • We classify existing methods into low-frequency and high-frequency matching based on their filter functions, highlighting their respective roles in encoding global colors and local textures.
@@
2 Preliminary
@@
Notations Let $\mathcal{T}=(H,Y)$ denote a real dataset, where $H$ represents the original data with $|H|=n$ samples, and $Y\in\mathbb{R}^{n\times c}$ is the one-hot label matrix for $c$ classes. The goal of DD is to learn a synthetic network $\mathcal{S}=(H_{s},Y_{s})$, where $|H_{s}|=m\ll n$ and $Y_{s}\in\mathbb{R}^{m\times c}$, such that models trained on $\mathcal{T}$ and $\mathcal{S}$ have comparable performance. In addition, there is a pre-trained distillation network $\phi(\cdot)$, such as ConvNet (DC) or ResNet-18 (ResNet) for the image datasets. We use $X=\phi(H)\in\mathbb{R}^{n\times d}$ and $X_{s}=\phi(H_{s})\in\mathbb{R}^{m\times d}$ to indicate the data representations learned by the distillation network on real and synthetic datasets, where $d$ is the dimension.
FFC and FLC matrices. For the real and synthetic datasets, the FFC matrices are defined as $X^{\top}X,X_{s}^{\top}X_{s}\in\mathbb{R}^{d\times d}$, respectively. These matrices capture correlations among feature channels and serve as important statistical descriptors of the datasets. Similarly, the FLC matrices are denoted as $X^{\top}Y,X_{s}^{\top}Y_{s}\in\mathbb{R}^{d\times c}$, respectively. Each column corresponds to the average representation of a class, capturing the alignment between channels and labels.
@@
Spectral Filtering. As $X^{\top}X$ is positive semi-definite, it can be decomposed into $X^{\top}X=U\Lambda U^{\top}$, where $U$ is the eigenvectors and $\Lambda$ is the diagonal eigenvalue matrix with $\lambda_{i}=\Lambda_{ii}$. The filtering function $f(\cdot)$ acting on $X^{\top}X$ is equivalent to acting on its eigenvalues, i.e., $f(X^{\top}X)=Uf(\Lambda)U^{\top}$. Without loss of generality, we assume that the eigenvalues are sorted in ascending order, i.e., $0\leq\lambda_{1}\leq\cdots\leq\lambda_{d}$. PCA points out that eigenvectors associated with large eigenvalues encode low-frequency information, while small eigenvalues correspond to high-frequency components. Therefore, a high-pass filter can be defined as $f(\lambda_{i})\geq f(\lambda_{i+1}),\forall i\in[1,d-1]$, as it will assign larger amplitudes to high-frequency eigenvectors, while the low-pass filters will do the opposite.
Spectral-domain Dataset Distillation. There are some DD methods operating in the spectral domain, such as FreD (FreD) and NSD (NSD). UniDD differs from them in two key aspects: First, FreD and NSD analyze the frequency of the input data itself, e.g., images in the Fourier domain, while UniDD operates on the spectral properties of the FFC matrix, making it more general and data-agnostic. Second, FreD and NSD remain data-parameterization methods whose optimization relies on existing objectives, while UniDD formulates a framework of DD objectives in the representation space, enabling the principled design of new distillation loss functions.
@@
Low-frequency DM $\left\|X^{\top}Y-X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=1$
DC $\left\|X^{\top}X-X_{s}^{\top}X_{s}\right\|_{F}^{2}+\left\|X^{\top}Y-X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=\{1,\lambda\}$
High-frequency MTT $\left\|(I-\alpha X^{\top}X)^{P}-(I-\alpha X_{s}^{\top}X_{s})^{Q}\right\|_{F}^{2}+$ $\alpha\left\|\sum_{p=0}^{P-1}(I-\alpha X^{\top}X)^{p}X^{\top}Y-\sum_{q=0}^{Q-1}(I-\alpha X_{s}^{\top}X_{s})^{q}X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=(1-\alpha\lambda)^{\{p,q\}}$
FRePo $\left\|(X^{\top}X+\beta I)^{-1}X^{\top}Y-(X_{s}^{\top}X_{s}+\beta I)^{-1}X_{s}^{\top}Y_{s}\right\|_{F}^{2}$ $f(\lambda)=(\lambda+\beta)^{-1}$
3 Spectral Filtering of Dataset Distillation
Before delving into the details, we first formally formulate UniDD, a spectral filtering framework that unifies diverse DD objectives.
Theorem 1.
There exists a wide range of dataset distillation methods that can be unified into a spectral filtering framework, defined as:
where the FFC and FLC matrices serve as a filter and signal, respectively, $f(\cdot)$ is the filtering function, and $g(\cdot)$ is a binary function with $g(X^{\top}Y)=I$ or $X^{\top}Y$.
We analyze four representative DD methods in the following and explain their relationship with different filter functions. Based on the filter behaviors, they can be divided into LFM-based and HFM-based methods. See Table 1 for a quick overview and Appendix B for more detailed derivations.
@@ 3.1 Low-frequency Matching
Statistical Matching. Matching important statistics between the real and synthetic datasets is a straightforward way to distill knowledge. DM (DM) proposes to match the average representations of each class, whose objective function can be defined as:
@@
where the corresponding filtering functions are $f(X^{\top}X)=I$ and $g(X^{\top}Y)=X^{\top}Y$, respectively. For clarity, we omit the mean normalization here, but this does not affect the definitions of $f$ and $g$.
However, DM does not perform well on complex distillation backbones, e.g., ResNet-18, as it only matches the statistic of the last layer of backbones. To overcome this limitation, SRe2(SRe2L) proposes to match the mean and variance information in each Batch Normalization (BN) layer. The objective function is defined as:
@@
Gradient Matching. Another LFM example is gradient matching (DC), which minimizes the differences of model gradients in the real and synthetic datasets. The gradients of the linear classifier are calculated as:
@@ 3.2 High-frequency Matching
Trajectory Matching. Matching the long-range training trajectories has been proven to be an effective approach for DD. We take MTT (MTT) as an example, which minimizes the differences between network parameters trained on the real and synthetic datasets. The objective function can be formulated as:
@@
where $\alpha$ is the learning rate and $W^{0}$ indicates the initial parameters. For $K$ iterations, we have:
@@
Subsequent works improve MTT from different perspectives, such as memory overhead (TESLA), trajectory training (FTD), and trajectory selection (DATM). However, they all follow the same objective function so that they can be naturally included in our unified framework.
KRR. The early-stage DD algorithms usually update the synthetic data by solving a two-level optimization problem. To reduce computation and memory costs, KRR-based methods are proposed to replace the inner-optimization with a closed-form solution:
@@
By applying a matrix identity transform (welling2013kernel) on the second term, the objective function can be unified into UniDD:
@@
It is worth noting that only the linear kernel case can be unified into our framework, e.g., KIP (KIP1; KIP2) and FrePo (FRePo). We leave the analysis of the non-linear kernel, e.g., Gaussian and polynomial, as further work.
4 In-depth Analysis of Filters
The above analysis reveals that DD methods adopt different filters, e.g., low-pass and high-pass. This section further justifies our claim and analyzes the roles of different filters through experiments on the Tiny-ImageNet dataset.
Power Spectral Density. We first verify that SM and GM behave as low-pass filters, whereas TM and KRR exhibit high-pass filtering. Specifically, Figure 1 visualizes the radially averaged power spectral density of the synthetic images learned by these four methods. The results show that SM and GM have more energy in the low-frequency band, while TM and KRR allocate greater power to high-frequency components, thereby supporting our claim.
Intra-class Similarity. We further investigate the influence of low-pass and high-pass filters on the synthetic data. Following G-VBSM (G-VBSM), we compute the intra-class cosine similarity of synthetic datasets distilled by different filters. As shown in Figure 3, the low-pass filter increases intra-class similarity of synthetic data and preserves the consistency of DD, while the high-pass filter improves data diversity by reducing intra-class similarity.
Visualization. Finally, we visualize the raw images and the synthetic images distilled by low-pass and high-pass filters. We observe that synthetic images produced by low-pass filters (Figure 3(b)) appear noticeably blurrier than the original images (Figure 3(a)), as low-pass filters assign higher weights to eigenvectors associated with larger eigenvalues. These eigenvectors represent the low-frequency components of the original data, e.g., coarse-grained shapes. In contrast, high-pass filters emphasize eigenvectors with smaller eigenvalues, thereby enhancing fine-grained and complex textures, as shown in Figure 3(c).
Figure 1: Power Spectral Density
Figure 2: Intra-class Similarity.
(a) Raw
(b) LP
(c) HP
Figure 3: Comparison between real and synthetic images. LP: Low-pass filter; HP: High-pass filter.